What is the answer to part e? (Please show your work)

The following is Randolph Corps contribution format income statement for last month: Sales Variable expenses Contribution maWhat is the answer to part e? (Please show your work)
Answer #1

Contribution margin ratio = 470000/1250000 = 37.60%

Break even sales = Fixed cost/Contribution margin ratio = 70000/.376 = 186170

Margin of safety in dollars = 1250000-186170 = 1063830